OpenStudy (anonymous):

A machine is 82% efficient. If it is able to do 1020 J of useful work, what will the total work or energy input have to be? Round your answer to the tens place. Be sure to include the correct unit in the answer.

OpenStudy (anonymous):

Can anyone help?

OpenStudy (anonymous):

If the machine is 82% efficient, what that means is that it turns 82% of the input energy into useful energy. To find out the input energy, think of this as a proportion. \[\frac{ 0.82 }{ 1020 }=\frac{ 1 }{ x }\rightarrow \frac{ 1020 }{ 0.82 }=x\] I don't have a calculator next to me, but that fraction should give you the correct answer.

OpenStudy (anonymous):

is it 836.4